Mary makes three invetments. The investment that pays 12% is twice the amount o the account that pays 9%. The amount invested at 10% is $500 more than the amount invested at 9%. If the annual interest income is $1555, how much money is invested at each rate?
Answer by checkley77(12844)

.12*2X+.09X+.10(X+500)=1,555
.24X+.09X+.10X+50=1,555
.43X=1,555-50
.43X=1,505
X=1,505/.43
X=3,500 AMOUNT INVESTED @ 9%
2*3,500=7,000 AMOUNT INVESTED @ 12%
3,500+500=4,000 AMOUNT INVESTED @ 10%
PROOF:
.12*7,000+.09*3,500+.10*4,000=1,555
840+315+400=1,555
1,555=1,555